# references.bib

### v4

@@ 5.3 Proof of the main theorem
First, we conclude the approach we used in the proof in section 5.1 as follows. We deformed the cosymbol $\sigma_{H}(P)$ to the operator $\sigma^{\prime}(P)(x,\xi)$ on $C_{0}(TM)$ by the Connes-Thom’s isomorphism. This allows both operators to reside on $TM$. Subsequently, a homotopy on symbol through rotation of coordinates on $T(TM)$ demonstrated that these two operators are homotopic.
Employing a similar approach for Theorem 3.2.6 allows us to give a proof for the main theorem (Theorem 5.2.2). We can try to transform the tangent-Clifford symbol into a symbol-like operator on $C_{0}(TM)$. Mainly, we need to prove that the two operators representing $KK$-classes $j^{G}([\sigma^{tcl}_{H}(P)]\otimes_{C^{\ast}(Gr(H)))}[\mathcal{E}])\otimes_{C^{\ast}(G,C_{0}(T^{\ast}M))}[\mathcal{D}_{M,\tilde{\Gamma}}^{cl}]$ and $[\mathcal{C}]\otimes_{C(M)}[P]$ are homotopic up to 00-elements in $KK$-classes.
@@
(of Theorem 5.2.2) Assume that $P_{0}$ is a transversally H-elliptic differential operator and $P$ is its bounded transformation. The proof is separated into the following five parts:
  • (i)
    Revert from $\mathbb{C}$ to $C(M)$ as in the proof of Theorem 3.2.6, the symbol class $[\sigma_{H}^{tcl}(P)]$ lies in $KK^{G}(C(M),C^{\ast}(Gr(H))\otimes_{C(M)}Cl_{\tilde{\Gamma}}(M)),$ and the index class $[P]$ lies in $KK(C^{\ast}(G,C(M)))$.
  • (ii)
    Take $\sigma^{\prime}(P)(x,\xi)$ to be the operator on $C_{0}(TM)$ we defined in Lemma 3.2.7. Then as in the proof of Theorem 3.2.6, we want to replace the class $[\sigma^{tcl}_{H}(P)]\otimes_{C^{\ast}(Gr(H)))}[\mathcal{E}]$ by some $[\sigma^{\prime}_{tcl}(P)]\in KK_{G}(C(M),C_{0}(T^{\ast}M)$. By definition, $\sigma^{tcl}_{H}(P)=\sigma_{H}(P)\hat{\otimes}1+(1-\sigma_{H}(P)^{2})^{1/2}\hat{\otimes}f_{\tilde{\Gamma}}$. The product over $C^{\ast}(Gr(H))$ with $[\mathcal{E}]$ transforms $\sigma_{H}(P)\hat{\otimes}1$ to $\sigma^{\prime}(P)\hat{\otimes}1$. Let $f_{\Gamma}$ is the ordinary cosymbol of the 00-order operator $\mathfrak{F}_{G}=\mathcal{D}_{G}(1+\mathcal{D}_{G}^{2})^{-1/2}$. Note that all the terms in $\mathcal{D}_{G}$ are normalized to 00-order, so the product with $[\mathcal{E}]$ transforms $f_{\tilde{\Gamma}}$ to $f_{{\Gamma}}$, and thus $(1-\sigma_{H}(P)^{2})^{1/2}\hat{\otimes}f_{\tilde{\Gamma}}$ to $(1-{\sigma^{\prime}(P)}^{2})^{1/2}\hat{\otimes}f_{{\Gamma}}$. The sum of the two resulting terms can be denoted by $\sigma^{\prime}_{tcl}(P)=\sigma^{\prime}(P)\hat{\otimes}1+(1-{\sigma^{\prime}(P)}^{2})^{1/2}\hat{\otimes}f_{{\Gamma}}$.
    Then we can represent the $KK$-product element $j^{G}([\sigma^{tcl}_{H}(P)])\otimes_{C^{\ast}(G,C^{\ast}(Gr(H)))}j^{G}([\mathcal{E}])$
    $\otimes_{C^{\ast}(G,C_{0}(T^{\ast}M))}[\mathcal{D}_{M,\tilde{\Gamma}}^{cl}]$ by the pairing
    $(L^{2}(E_{1}\hat{\otimes}_{Cl_{\Gamma}(M)}L^{2}(\Lambda^{\ast}(M)),Q=N_{1}(\sigma_{tcl}^{\prime}(P)\hat{\otimes}1)+N_{2}(1\otimes\widetilde{\mathcal{D}_{M,\tilde{\Gamma}}^{cl}})),$
    where $E_{1}$ is the Hilbert module of the sections of the family of Hilbert spaces $\{E_{x}\hat{\otimes}Cl_{\tilde{\Gamma}_{x}}(M)\otimes L^{2}(T_{x}M)\}_{x\in M}$, $N_{1},N_{2}$ are pseudo-differential operators constructed in the $KK$-product, and $\widetilde{\mathcal{D}_{M,\tilde{\Gamma}}^{cl}}$ is the $KK$-theoretic connection for $\mathcal{D}_{M,\tilde{\Gamma}}^{cl}$. Note that the Clifford terms act over the space $\Lambda^{\ast}(TM)$.
  • (iii)
    Let $\xi$ be the multiplication variable on $TM$, $\chi$ and $\zeta$ be the derivation variables. The operator $\sigma^{\prime}(P)$ and the symbol term $f_{\Gamma}$ take variables in $x$ and $\xi$, and the symbol of $\mathcal{D}_{M,\tilde{\Gamma}}^{cl}$ takes variables in $\chi$ and $\zeta$. Our rotation homotopy is defined using the symbol of $\sigma^{\prime}(P)\otimes 1$ and $\widetilde{\mathcal{D}^{cl}_{M,\tilde{\Gamma}}}$ by rotation of variables
    $\xi\mapsto\cos t\ \xi-\sin t\ \chi;\ \chi\mapsto\sin t\ \xi+\cos t\ \chi,\ t\in[0,\pi/2].$
    Note that $f_{\Gamma}$ depends on the orbits whose dimensions may jump at some points, so we construct the rotation homotopy to be induced from the rotation homotopy of the $\mathfrak{b}$ symbol of the orbital Dirac operator $\mathfrak{F}_{G}$, defined in Definition 5.5 in [kasparov2022k], on the trivial vector bundle $\mathfrak{g}_{M}\cong M\times\mathfrak{g}$. The variables corresponding to $\xi,\chi$ for the $\mathfrak{b}$ symbol are $\eta$ and $\theta$. The rotation homotopy for $f_{\Gamma}$ is taken between $\eta$ and $\theta$.
  • (iv)
    Such a homotopy involves both the multiplication variable and derivation variable, so it is necessary to check that the commutator of each two terms is 00 in their corresponding $KK$-classes, so that the homotopy can be well-defined in the operator sense. The only non-trivial case we need to consider is that $f_{\Gamma}$ does not commute well with the operator $\widetilde{\mathcal{D}^{cl}_{M,\tilde{\Gamma}}}$ because these two operators use the same Clifford variables. However, the algebra $C^{\ast}(G,C(M))$ contains the multiplier $(1+\Delta_{G})^{-1}$ which kills the non-trivial commutators. Along this rotation homotopy, the commutator $[f_{\Gamma},\widetilde{\mathcal{D}^{cl}_{M,\tilde{\Gamma}}}]$ will not be killed by the multiplication with $C^{\ast}(G,C(M))$. However, by the important remark in the proof of Theorem 10.3 of [kasparov2022k], the homotopy can be corrected to have norm $1$ and at the end of the homotopy, the non-trivial commutator again vanishes. As a result, we get a well-defined rotation homotopy as desired.
  • (v)
@@
    Use functional calculus on the operator $P$ such that $\lVert P\rVert\leq 1$. Consider the operator $T=P\hat{\otimes}1+(1-{P}^{2})^{1/2}\hat{\otimes}\mathfrak{F}_{G}$ corresponding to the symbol class $[\sigma^{tcl}_{H}(P)]\otimes_{C^{\ast}(Gr(H))}[\mathcal{E}]$ act on $L^{2}(E_{1}\hat{\otimes}\Lambda^{\ast}(TM))$. It suffices to show that $1-T^{2}$ is a compact operator, so that such an operator $T$ provides a representation of a $KK$-class. As in the proof of Theorem 10.3 of [kasparov2022k], the commutator term $[P\hat{\otimes}1,1\hat{\otimes}\mathfrak{F}_{G}]$ is estimated on two sides by the operator $(1+\mathcal{D}_{O}^{2})^{-1/2}$, i.e., for some positive constant number $c$, we have
    $-c(1+\mathcal{D}_{O}^{2})^{-1/2}\leq[P\hat{\otimes}1,1\hat{\otimes}\mathfrak{F}_{G}]\leq c(1+\mathcal{D}_{O}^{2})^{-1/2}.$
@@
    By calculation, $1-T^{2}=(1-P^{2})(1-\mathfrak{F}_{G}^{2})-(1-P^{2})^{1/2}[P\hat{\otimes}1,1\hat{\otimes}\mathfrak{F}_{G}]$. Note that by definition, the transversally H-elliptic operator $P$ satisfies that $(1-\sigma_{H}(P)^{2})(1+\sigma_{H}(\Delta_{G}))^{-1}$ is compact and thus has negative H-order. So by Theorem 1.6.13, $(1-P^{2})(1+\mathcal{D}_{O}^{2})^{-1}$ is a compact operator. Thus $1-T^{2}$ is also compact.
    The operator $\sigma^{\prime}(P)(x,\xi)$ is defined on $TM$ and the operator $P$ can be lifted to $TM$ with the same symbol, to be a $KK$-theoretic connection $\tilde{P}$. Note that $\sigma^{\prime}(P_{0})(x,\xi)$ is the Fourier transform of the freezing coefficient of the H-principal part of $P_{0}$, so the rotation homotopy transforms the multiplication with $\xi$-variable by the differential by $x$-variable in the operator sense. $\sigma^{\prime}(P_{0})(x,\xi)$ on $TM$ is a polynomial in $\xi$ variable, so it is transformed to an operator having the same $H$-principal part with $\tilde{P_{0}}$ on $TM$. Thus by functional calculus, $\sigma^{\prime}(P)(x,\xi)$ on $TM$ is transformed to $\tilde{P}$ up to lower order terms. As a result, we see that the lift of $T=P\hat{\otimes}1+(1-{P}^{2})^{1/2}\hat{\otimes}\mathfrak{F}_{G}$ to $TM$ gives a representation of the operator corresponding to the symbol $\sigma^{\prime}(P)(x,\chi)\hat{\otimes}1+(1-\sigma^{\prime}(P)^{2}(x,\chi))^{1/2}\hat{\otimes}f_{{\Gamma}}(x,\chi)$ after the rotation homotopy on $\sigma^{\prime}_{tcl}(P)$.
    As in the proof of Theorem 3.2.6, by executing the rotation homotopy, the Dolbeault component $\mathcal{D}_{M,\tilde{\Gamma}}^{cl}$ in $Q$ is transitioned into the assembly of Bott–Dirac operators $\mathcal{C}=\{\mathcal{C}_{x},x\in M\}$ on $L^{2}(\Lambda^{\ast}(T_{x}M))$.
    By the statements above, $Q$ can be transformed by the rotation homotopy into the operator $N_{1}\tilde{T}+N_{2}1\hat{\otimes}\mathcal{C}$, representing the $KK$-product of $[T]$ and $[\mathcal{C}]$. The operator $\mathfrak{F}_{G}$ in $T$ can be removed since it is killed by the multiplication with $C(M)\cdot(1+\Delta_{G})^{-1}\subset C^{\ast}(G,C(M))$. Note also that the Bott-Dirac operators $\mathcal{C}$ correspond to the $1$ element in $KK(C^{\ast}(G,C(M)),C^{\ast}(G,C(M)))$. As a result, after the rotation, we arrive at the index class $[P]\in KK(C^{\ast}(G,C(M)))$.
  • (vi)
    Consider the induced restriction from $KK^{0}(C^{\ast}(G,C(M)))$ to $KK^{0}(C^{\ast}(G))$ and from $KK^{G}(C(M),C^{\ast}(Gr(H))\otimes_{C(M)}Cl_{\tilde{\Gamma}}(M))$ to $KK^{G}(\mathbb{C},C^{\ast}(Gr(H))\otimes_{C(M)}Cl_{\tilde{\Gamma}}(M))$ by the inclusion $\mathbb{C}\subset C(M)$, we will get analytical index class $[P]$ which is the left-hand side.
@@
It is considerable to establish the index theory for transversally H-elliptic operators via an adiabatic groupoid approach, similar to the construction in Section 3.1. For each representation of $G$, it entails contemplating distinct $C^{\ast}$-algebras for each fiber of the adiabatic groupoid, so that it can represent a commutative diagram and we obtain an index in $\mathbb{Z}$. By such construction, the $C^{\ast}$-algebra diagram induces an element in $\text{Hom}(R(G),\mathbb{Z})\cong\widehat{R(G)}\cong K^{0}(C^{\ast}(G))$, and it should be consistent with the analytical index class of $P$. A different proof of the index theorem using such a groupoid approach is left for future research.

### v5

@@ 5.3 Proof of the main theorem
(of Theorem 5.1.4) Assume that $P$ is a $G$-invariant transversally H-elliptic differential operator of order 00. The proof can be separated into the following steps:
  1. (i)
    We first verify each element of the $KK$-product on the right-hand side.
    • The tangent-Clifford cosymbol element is $[\sigma^{tcl}_{H}(P)]\in KK^{G}(C_{0}(M),C_{0}(M;E)\hat{\otimes}_{C_{0}(M)}$ $C^{\ast}(Gr(H))\hat{\otimes}_{C_{0}(M)}Cl_{\Gamma}(M))$. The Hilbert module is the family $\{E_{x}\hat{\otimes}C^{\ast}(Gr(H)_{x})$$\hat{\otimes}Cl_{\Gamma_{x}}(M)\}_{x\in M}$, and the operator is the family of operators $\{\sigma^{tcl}_{H}(P,x)\}_{x\in M}$ we defined in Definition 4.5.4.
    • $[\eta_{C^{\ast}(Gr(H))}]\otimes_{C_{0}(M)}1_{Cl_{\Gamma}(M)}\in\mathcal{R}KK(M;C^{\ast}(Gr(H))\hat{\otimes}_{C_{0}(M)}Cl_{\Gamma}(M),Cl_{\tau\oplus\Gamma}(M))$. The Hilbert module is the family $\{L^{2}(Gr(H)_{x})\hat{\otimes}Cl_{\tau_{x}\oplus\Gamma_{x}}(M)\}_{x\in M}$, and for each $x$ the operator is the covector fields $\eta_{x}(y)$ over $Gr(H)_{x}$.
    • $[d_{M,\Gamma}]\in K^{0}(C^{\ast}(G,Cl_{\tau\oplus\Gamma}(M)))$. The Hilbert space is $H=L^{2}(\Lambda^{\ast}(M))$ with $Cl_{\tau\oplus\Gamma}(M)$ action by Clifford multiplication operator on covectors by
      $\xi_{1}\oplus\xi_{2}\mapsto\text{ext}(\xi_{1})+\text{int}(\xi_{1})+i(\text{ext}(\xi_{2})-\text{int}(\xi_{2})),$
      and the operator is $F_{M}=D_{M}/(1+D_{M}^{2})^{1/2}$, where $D_{M}=d_{M}+d_{M}^{\ast}$ and $d_{M}$ is the operator of exterior derivation on $H$.
    The $KK$-product on the right-hand side can be represented on the Hilbert space =$\mathcal{H}=\{E_{x}\hat{\otimes}L^{2}(Gr(H)_{x})\hat{\otimes}Cl_{\tau_{x}\oplus\Gamma_{x}}(M)\}_{x\in M}\hat{\otimes}_{Cl_{\tau\oplus\Gamma}(M)}L^{2}(\Lambda^{\ast}(M))$ with a $C^{\ast}(G,C_{0}(M))$ action. The tangent-Clifford cosymbol $\sigma^{tcl}_{H}(P,x)$ acts on $\{E_{x}\hat{\otimes}$ $L^{2}(Gr(H)_{x})\hat{\otimes}Cl_{\Gamma_{x}}(M)\}_{x\in M}$, the covector field $\eta_{x}$ acts on $\{L^{2}(Gr(H)_{x})\hat{\otimes}Cl_{\tau_{x}\oplus\Gamma_{x}}(M)\}_{x\in M}$$\hat{\otimes}_{Cl_{\tau\oplus\Gamma}(M)}L^{2}(\Lambda^{\ast}(M))$, and $F_{M}$ acts on $L^{2}(\Lambda^{\ast}(M))$.
    The triple $KK$-product can be represented as the operator on the Hilbert space =$\mathcal{H}=\{E_{x}\hat{\otimes}L^{2}(Gr(H)_{x})\hat{\otimes}Cl_{\tau_{x}\oplus\Gamma_{x}}(M)\}\hat{\otimes}_{Cl_{\tau\oplus\Gamma}(M)}L^{2}(\Lambda^{\ast}(M))$:
    $Q=1\hat{\otimes}\eta\hat{\otimes}1+(1-\eta^{2})^{1/2}(N_{1}\sigma_{H}^{tcl}(P)\hat{\otimes}1\hat{\otimes}1+N_{2}1\hat{\otimes}1\hat{\otimes}F_{M}),$
    where $\eta=\{\eta_{x}\}_{x\in M}$ and $\sigma^{tcl}_{H}(P)=\{\sigma^{tcl}_{H}(P,x)\}_{x\in M}$, with appropriate operators $N_{1}$ and $N_{2}$ such that $N_{1}^{2}+N_{2}^{2}=1$.
  2. (ii)
    We can choose a small radical neighborhood $\mathcal{V}_{x}$ of $x$ with the radii $r_{x}$ varying smoothly for $x\in M$ as in Subsection 5.2.1 and 5.2.2. Assume that $\lVert\eta_{x}\rVert=1$ outside of $\mathcal{V}_{x}$ (or a smaller neighborhood) by renormalization. By this renormalization of $\eta_{x}$, the product $Q$ only depends on the value of $\eta_{x}$ and $\sigma^{tcl}_{H}(P,x)$ on $\mathcal{V}_{x}\subset Gr(H)_{x}$. By the local homeomorphism between $\mathcal{V}_{x}\subset Gr(H)_{x}$ and $U_{x}\subset M$, $\sigma^{tcl}_{H}(P,x)$ will be transformed into the operator $T^{x}$ with constant cosymbol $\sigma^{tcl}_{H}(P,x)$ on $U_{x}$, and $\eta_{x}(y)$, which is a covector field over $\mathcal{V}_{x}\subset Gr(H)_{x}$, will be transformed into a covector field on $U_{x}$. We still denote it by $\eta_{x}(y)$ for $y\in U_{x}\subset M$. As a result, we can transform all these operators to a small neighborhood $U$ of the diagonal $M\times M$ with $U_{x}=U\cap(M\times\{x\})$.
    Then after the local homeomorphism, the operator $\sigma_{H}(P,x)$ and $\eta_{x}$ are defined on $E_{x}\hat{\otimes}L^{2}(Gr(H)_{x})$ will be transformed into operators on $E_{x}\hat{\otimes}L^{2}(U_{x})\cong L^{2}(E|_{U_{x}})$ by the trivialization condition of $E$ on $U_{x}$. So the whole operator $Q$ is transformed into
    $Q^{\prime}=1\hat{\otimes}\eta\hat{\otimes}1+(1-\eta^{2})^{1/2}(N_{1}\{T^{x}\}_{x\in M}\hat{\otimes}1\hat{\otimes}1+N_{2}1\hat{\otimes}1\hat{\otimes}F_{M}),$
    acting on the Hilbert space =$\mathcal{H}^{\prime}=\{L^{2}(E|_{U_{x}})\hat{\otimes}Cl_{\tau_{x}\oplus\Gamma_{x}}(M)\}_{x\in M}\hat{\otimes}_{Cl_{\tau\oplus\Gamma}(M)}L^{2}(\Lambda^{\ast}(M))=\{L^{2}(E)$$\hat{\otimes}L^{2}(\Lambda^{\ast}(U_{x}))\}_{x\in M}$. Consequently, $Q^{\prime}$ can be considered as a family of operators $Q^{\prime}_{x}$ on $\mathcal{H}^{\prime}_{x}=L^{2}(E)\hat{\otimes}L^{2}(\Lambda^{\ast}(U_{x}))$, parametrized by $x\in M$:
    $Q^{\prime}_{x}=1\hat{\otimes}\eta_{x}\hat{\otimes}1+(1-\eta_{x}^{2})^{1/2}(N_{1}T^{x}\hat{\otimes}1\hat{\otimes}1+N_{2}1\hat{\otimes}1\hat{\otimes}F_{M}).$
    The operator $T^{x}$ can be represented as
    $T^{x}=P(x)\hat{\otimes}1+(1-P(x)^{2})^{1/2}\hat{\otimes}\mathfrak{F}_{G}(x),$
[…]
    for all $y\in U_{x}$. By the arguments in Section 5.3 in [kasparov2022k], the function from orbit $\mathcal{O}$ to the operator $\mathfrak{F}_{\mathcal{O}}=\mathcal{D}_{\mathcal{O}}/(1+\mathcal{D}^{2}_{\mathcal{O}})^{1/2}$ defined in Section 4.2 is continuous in the $\ast$-strong topology, and $1-\mathfrak{F}_{\mathcal{O}}^{2}$ is continuous in norm, so the above homotopy provides a standard homotopy between $KK$-cycles, from $T^{x}$ to $T=P\hat{\otimes}1+(1-P^{2})^{1/2}\hat{\otimes}\mathfrak{F}_{G}$ on $U_{x}$. During the homotopy, we shall note that these operators are considered on the consistent Hilbert space $L^{2}(E)$.
@@
    It suffices to show that $a(1-T^{2})$ is a compact operator for arbitrary $a\in C_{0}(M)$, so that such an operator $T$ actually represents a $KK$-cycle. As in the proof of Theorem 10.3 of [kasparov2022k], the commutator term $[P\hat{\otimes}1,1\hat{\otimes}\mathfrak{F}_{G}]$ is estimated on two sides by the operator $(1+\mathcal{D}_{\mathcal{O}}^{2})^{-1/2}$, i.e., for some positive constant number $c$, we have
    $-c(1+\mathcal{D}_{\mathcal{O}}^{2})^{-1/2}\leq[P\hat{\otimes}1,1\hat{\otimes}\mathfrak{F}_{G}]\leq c(1+\mathcal{D}_{\mathcal{O}}^{2})^{-1/2}.$
@@
    By calculation, $1-T^{2}=(1-P^{2})(1-\mathfrak{F}_{G}^{2})-(1-P^{2})^{1/2}[P\hat{\otimes}1,1\hat{\otimes}\mathfrak{F}_{G}]$. Note that by definition, the transversally H-elliptic operator $P$ satisfies that $(1-\sigma_{H}(P)^{2})(1+\sigma_{H}(\Delta_{G}))^{-1}$ is compact and thus has negative H-order. So by Remark 1.6.14, $(1-P^{2})(1+\mathcal{D}_{\mathcal{O}}^{2})^{-1}$ also has negative order. Thus $a(1-T^{2})$ is compact for arbitrary $a\in C_{0}(M)$ acting on the first copy $M$ of $U\subset M\times M$ where $T$ lies.
[…]
    $\phi_{t}:a\mapsto p_{t}^{\ast}(a)\in C_{b}(M\times M)$
    gives a homotopy between the homomorphism of multiplication on $C_{b}(U)$ by $C_{0}(M)$ on the first and second $M$ variable.
[…]
    Then, by the homotopies constructed above, $Q^{\prime}$ is transformed into a family of operators $Q^{\prime\prime}=\{Q_{x}^{\prime\prime}\}_{x\in M}$:
    $Q^{\prime\prime}_{x}=1\hat{\otimes}\Theta_{x}\hat{\otimes}1+(1-\Theta_{x}^{2})^{1/2}(N_{1}P\hat{\otimes}1\hat{\otimes}1+N_{2}1\hat{\otimes}1\hat{\otimes}F_{M}),$
    acting on the Hilbert space $\mathcal{H}^{\prime}_{x}=L^{2}(E)\hat{\otimes}L^{2}(\Lambda^{\ast}(U_{x}))$ for each $x\in M$.
  3. (iii)
    We calculate the $KK$-product $[\Theta_{M}]\otimes_{Cl_{\tau}(M)}[d_{M}]\otimes_{C_{0}(M)}[P]$, in which $[P]$ takes the product with $[\Theta_{M}]\in KK(C_{0}(M),C_{0}(M)\hat{\otimes}Cl_{\tau}(M))$ over $C_{0}(M)$ in the second variable and $[d_{M}]$ takes the product with $[\Theta_{M}]$ over $Cl_{\tau}(M)$. First, we verify each element of this product.
    • The local dual Dirac element $[\Theta_{M}]\in KK(C_{0}(M),C_{0}(M)\hat{\otimes}Cl_{\tau}(M))$ is defined by the field of pairs $\{(Cl_{\tau}(U_{x}),\Theta_{x})_{x\in M}\}$.
    • The index class $[P]\in K^{0}(C^{\ast}(G,C_{0}(M)))$ is defined by the pair $(L^{2}(E),P)$.
    • The Dirac element $[d_{M}]\in K^{0}(Cl_{\tau}(M))$ is defined by the pair $(L^{2}(\Lambda^{\ast}(M)),F_{M}=D_{M}/(1+D_{M}^{2})^{1/2})$, where $D_{M}=d_{M}+d_{M}^{\ast}$.
    So the triple $KK$-product $[\Theta_{M}]\otimes_{Cl_{\tau}(M)}[d_{M}]\otimes_{C_{0}(M)}[P]$ can be written as the pair
    $(C^{\ast}(G,C_{0}(M))\to\mathcal{H}^{\prime},\mathcal{Q}),$
    where $\mathcal{H}^{\prime}$ is the Hilbert space $\{L^{2}(E)\hat{\otimes}L^{2}(\Lambda^{\ast}(U_{x}))\}_{x\in M}$, which can be considered as a family of Hilbert spaces $\mathcal{H}^{\prime}_{x}=L^{2}(E)\hat{\otimes}L^{2}(\Lambda^{\ast}(U_{x})$, and the operator $\mathcal{Q}=\{\mathcal{Q}_{x}\}_{x\in M}$ is a family of operators $\mathcal{Q}_{x}$ on $\mathcal{H}^{\prime}_{x}$ parametrized by $x\in M$:
    $\mathcal{Q}_{x}=N_{1}P\hat{\otimes}1\hat{\otimes}1+N_{2}[1\hat{\otimes}\Theta_{x}\hat{\otimes}1+(1-\Theta_{x}^{2})^{1/2}1\hat{\otimes}1\hat{\otimes}F_{M}]$
    with $N_{1}^{2}+N_{2}^{2}=1$.
    By the construction of $KK$-product (Theorem 2.11 in [kasparov1988equivariant]), the operators $N_{1}$ and $N_{2}$ are assigned to commute with $C_{0}(M)$, $P\hat{\otimes}1\hat{\otimes}1$, $11\hat{\otimes}\Theta_{x}\hat{\otimes}$ 1 and $1\hat{\otimes}1\hat{\otimes}F_{M}$ module $\mathcal{K}(L^{2}(M\times M))$, and they satisfy $N_{1}(\mathcal{K}(L^{2}(M))\otimes 1)\subset\mathcal{K}(L^{2}(M\times M))$ so $N_{1}$ vanishes $[P\hat{\otimes}1\hat{\otimes}1,1\hat{\otimes}\Theta_{x}\hat{\otimes}1]$ and $1-\Theta_{x}^{2}$, $N_{2}(1\otimes\mathcal{K}(L^{2}(M))\subset\mathcal{K}(L^{2}(M\times M))$ and $N_{2}$ vanishes $[1\hat{\otimes}\Theta_{x}\hat{\otimes}1,1\hat{\otimes}1\hat{\otimes}F_{M}]$ and $[P\hat{\otimes}1\hat{\otimes}1,1\hat{\otimes}1\hat{\otimes}F_{M}]$. So we see that
    $a[Q^{\prime\prime},\mathcal{Q}]a^{\ast}\geq 0\ \text{mod}\ \mathcal{K}(\mathcal{H}^{\prime})$
    for all $a\in C^{\ast}(G,C_{0}(M))$. Then by Lemma 11 in [skandalis1984some], the $KK$-cycles $(\mathcal{H}^{\prime},Q^{\prime\prime})$ and $(\mathcal{H}^{\prime},\mathcal{Q})$ are operator homotopic. As a result, we obtain the following equation:
    $[\Theta_{M}]\otimes_{Cl_{\tau}(M)}[d_{M}]\otimes_{C_{0}(M)}[P]$
    $=$ $j^{G}([\sigma^{tcl}_{H}(P)])\otimes_{C^{\ast}(G,C^{\ast}(Gr(H))\hat{\otimes}_{C_{0}(M)}Cl_{\Gamma}(M))}[\mathcal{D}_{M,{\Gamma}}^{cl^{\prime}}]\in K^{0}(C^{\ast}(G,C_{0}(M))).$
  4. (iv)
    By Theorem 4.8 in [kasparov1988equivariant],
    $[\Theta_{M}]\otimes_{Cl_{\tau}(M)}[d_{M}]=[1]\in KK(C_{0}(M),C_{0}(M)),$
    so its product with index class $[P]\in K^{0}(C^{\ast}(G,C_{0}(M)))$ over $C_{0}(M)$ will be $[P]$ itself. So we obtain the index class $[P]\in K^{0}(C^{\ast}(G,C_{0}(M)))$ on the left-hand side as we desired.
@@
It is also considerable to establish the index theory for transversally H-elliptic operators via an adiabatic groupoid approach, similar to the construction in Section 3.1. For each representation of $G$, it entails contemplating distinct $C^{\ast}$-algebras for each fiber of the adiabatic groupoid, so that it can represent a commutative diagram and we obtain an index in $\mathbb{Z}$. By such construction, the $C^{\ast}$-algebra diagram induces an element in $\text{Hom}(R(G),\mathbb{Z})\cong\widehat{R(G)}\cong K^{0}(C^{\ast}(G))$, and it should be consistent with the distributional index class of $P$ we discussed in Remark 4.5.8. A different proof of the index theorem using such a groupoid approach is left for future research.
@@
In this chapter, we focus on verifying the compactness condition outlined in Definition 4.5.1, i.e.,
Question 6.0.1.
Let $e$ be an fixed element in the multiplier algebra of the $C^{\ast}(Gr(H)_{x})$-module $E=C^{\ast}(Gr(H)_{x})$, $e\in\mathcal{M}(C^{\ast}(Gr(H)_{x}))$, how to show whether or not $e$ lies in the compact operator space $\mathcal{K}(E)=E$.